A father is 24 years older than is son in 7 years the son age will be 2/5 of his father. What are their present age.

Answer by mananth(16946)

A father is 24 years older than is son in 7 years the son age will be 2/5 of his father. What are their present age.
son =x years
father = x+24
...
in 7 years
son = x+7
father = x+24+7 = x+31
...
x+7 = 2/5 (x+31)
5(x+7)=2(x+31)
5x+35=2x+62
5x-2x=62-35
3x=27
/3
x= 9 the son's age
father = x+24 = 33 years
...
m.ananth@hotmail.ca